# On the nullspace of split graphs

### v1

@@ Abstract
We study the nullspace of the adjacency matrix of split graphs, whose vertex set can be partitioned into a clique and an independent set. We introduce the clique-kernel, a subspace that determines when vertices of the clique appear in the support of a kernel eigenvector. We prove that this subspace is at most one-dimensional, leading to the formula
$\operatorname{nl}(Sp)=\operatorname{nl}(R)+\dim(\mathrm{Cker}(Sp)),$
which fully describes the nullity of a split graph in terms of the bipartite adjacency submatrix $R$. We also analyze unbalanced split graphs through the concept of swing vertices and characterize the structure of their kernel supports. Furthermore, we study the behavior of the nullspace under Tyshkevich composition and derive a closed formula for the determinant using the Schur complement. These results provide a unified algebraic framework for understanding when a split graph is singular and how its combinatorial structure determines its nullspace.
@@ 1 Introduction
The study of spectral properties of graphs, particularly the nullspace of the adjacency matrix, has emerged as a fundamental area in algebraic graph theory with significant implications for both theoretical mathematics and applied sciences. A graph $G$ is said to be singular if its adjacency matrix $A(G)$ is singular, that is, if the dimension of the nullspace $\operatorname{\mathcal{N}}(A(G))$ is at least one. The nullity $\operatorname{nl}(G)$ of a graph $G$ is the multiplicity of zero as an eigenvalue of $A(G)$, equivalently, the dimension of $\operatorname{\mathcal{N}}(A(G))$. Understanding the structure of singular graphs and characterizing their nullspace has been a subject of intensive investigation since 1957 when von Collatz and Sinogowitz posed the problem, see [von1957spektren].
@@
Sciriha’s seminal work [sciriha2007characterization] established fundamental concepts for studying singular graphs through the introduction of core vertices—those vertices corresponding to non-zero entries in kernel eigenvectors. This approach led to a systematic characterization of singular graphs in terms of minimal configurations and core structures. Subsequently, numerous researchers have contributed to this framework, including studies on nut graphs (core graphs of nullity one) [sciriha1998construction], maximal core sizes [sciriha2009maximal], and the relationship between core and Fiedler vertices [ali2016coalescing]. Jaume and Molina, see [jaume2018null], presented a decomposition of trees in terms of the supported vertices, which turn out to be equivalent to the classical Edmond-Gallai decomposition, see [jaume2021null].
@@
Graphs whose vertex set can be partitioned into a clique and an independent set were first singled out by Földes and Hammer in 1977, under the name split graphs. The same authors had earlier studied them implicitly as the graphs that do not contain an induced copy of $2K_{2}$, $C_{4}$ or $C_{5}$ The term “split” was coined to stress that the vertex set is split into two sets whose induced subgraphs are as dissimilar as possible. Subsequent monographs of Golumbic [golumbic2004algorithmic] and Mahadev–Peled [mahadev1995threshold] embedded the class into the wider hierarchy of perfect graphs and made the notion standard in graph theory. Split graphs admit efficient algorithms for many NP-hard problems that are intractable on general graphs [hammer1981splittance]. Moreover, split graphs can be recognized in linear time, and classical problems such as finding maximum cliques and maximum independent sets can be solved efficiently on this class [foldes1977split].
@@
The spectral properties of split graphs have received considerable attention. Threshold graphs, which form a proper subclass of split graphs characterized by the absence of induced $P_{4}$, $C_{4}$, and $2K_{2}$ [mahadev1995threshold], have been extensively studied from a spectral perspective. Sciriha [sciriha2011spectrum] investigated the spectrum of threshold graphs and established that singular threshold graphs possess kernel eigenvectors with at most two non-zero entries. Recent work by Panelo, Jaume, and Machado Toledo [panelo5316214core] on the core-nilpotent decomposition of threshold graphs further illuminates the algebraic structure of these important graph families.
@@
Despite the rich literature on both singular graphs and split graphs, a comprehensive study of the nullspace structure specifically for split graphs has been lacking. The block structure of the adjacency matrix of a split graph,
$A(S)=\begin{pmatrix}J-I&R\\ R^{t}&0\end{pmatrix}$,
@@
where $J$ is the all-ones matrix, $I$ is the identity matrix, and $R$ encodes adjacencies between the clique $K$ and independent set $I$, naturally suggests a systematic approach to understanding $\ker(Sp)$. This structure allows us to decompose the nullspace problem into more manageable components and to establish clear relationships between the rank properties of $R$ and the nullity of $S$.
@@
    A detailed analysis of unbalanced split graphs (Section 4)—those graphs where $|K|\neq\omega(Sp)$ or $|S|\neq\alpha(Sp)$. We introduce the concept of swing vertices, which can belong to either part of a bipartition, and prove that the nullspace structure depends critically on whether the set of swing vertices forms a clique or an independent set.
@@
    For split graphs with supported clique vertices (Section 5), we prove that the nullity equals $\operatorname{nl}(R)+1$ when the support intersects the clique non-trivially (Theorem 5.3), and we provide explicit formulas relating kernel eigenvectors to the adjugate matrix when $\operatorname{nl}(Sp)=1$ (Theorem 5.7).
@@
    An explicit determinant formula for split graphs (Section 7) expressed in terms of the matrix $R$ using the Schur complement and the matrix determinant lemma (Theorem 7.3), which facilitates computational verification of singularity.
@@
    Applications to the Tyshkevich composition (Section 6) and the intersection graph construction (Section 8), demonstrating how our nullspace characterizations extend to composite structures.
Our results extend and complement existing work on singular graphs by exploiting the special structure of split graphs. While Sciriha’s general theory [sciriha2007characterization] applies to arbitrary graphs, the block decomposition of split graph adjacency matrices enables us to obtain more precise characterizations. Furthermore, our treatment of unbalanced split graphs and swing vertices reveals phenomena that have not been systematically explored in the literature on singular graphs.
The remainder of this paper is organized as follows. Section 2 introduces notation and provides background on split graphs. Section 3 establishes general facts about the kernel of split graphs, including fundamental relationships between $\operatorname{\mathcal{N}}(R)$, $\operatorname{\mathcal{N}}(R^{t})$, and $\operatorname{\mathcal{N}}(Sp)$. Section 4 analyzes unbalanced split graphs and characterizes swing vertices. Section 5 focuses on split graphs whose kernel eigenvectors have support intersecting the clique. Section 6 studies the Tyshkevich composition. Section 7 derives an explicit determinant formula. Section 8 introduces the intersection graph construction and establishes conditions under which $\text{null}(Sp)\leq 1$.
Throughout this work, we demonstrate that the nullspace of split graphs exhibits rich structural properties that can be understood through the interplay between the clique-independent set bipartition and the rank properties of the adjacency submatrix $R$. Our results provide both theoretical insights into the spectral theory of structured graph classes and computational tools for determining singularity and nullity of split graphs.
We believe that the present work fills a natural gap in the spectral theory of split graphs and offers a concise, matrix-free toolkit that may be useful in both theoretical and algorithmic contexts.
@@ 2 Notation
If $A$ is a matrix, then $A_{ij}$ denotes the $(i,j)$-entry of $A$ (i.e., the entry of $A$ in the intersection of row $i$ and column $j$), $A_{i*}$ denotes the row $i$ of $A$, and $A_{*j}$ denotes the column $j$ of $A$. We denote by $A(G)$ the adjacency matrix of $G$. If $G$ is clear from the context we just write $A$. By $I$ we denote the identity matrix. The range of $A$, i.e., the vector space spanned by the columns of $A$, is denoted by $\operatorname{\mathcal{R}}(A)$. Its dimension, the rank of $A$, is denoted by $\operatorname{rk}(A)$. The null space of $A$ is denoted by $\operatorname{\mathcal{N}}(A)$. Its dimension, the nullity of $A$, is denoted by $\operatorname{nl}(A)$. If $A$ is square, we denote the adjugate of $A$ by $\operatorname{adj}(A)$.
Let $U$ be a set of vector of some vector space $V$, by $\langle U\rangle$ we denote the subspace spanned by all the linear combinations of vectors in $U$. We just write $\langle u\rangle$ instead of $\langle\{u\}\rangle$.
$J$ is the “all-ones” matrix. Since $\operatorname{rk}(J)=1$, $\operatorname{nl}(J)$ is the number of columns of $J$ minus 1. $\operatorname{1\!\!1}=(1,\ldots,1)^{t}$ is the “all-ones” vector (i.e., the matrix $J$ with just one column). If $A$ is a non-zero matrix with non-negative entries, note that $\operatorname{1\!\!1}\notin\operatorname{\mathcal{N}}(A)$. Clearly, $\operatorname{\mathcal{R}}(J)=\langle\operatorname{1\!\!1}\rangle$.
If $x=(x_{1},\ldots,x_{n})^{t}\in\mathbb{R}^{n}$, where “t” means transpose, we have $\operatorname{1\!\!1}^{t}x=\sum_{i=1}^{n}x_{i}$. Clearly, $\operatorname{1\!\!1}^{t}x=0$ if and only if $x\in\operatorname{\mathcal{N}}(J)$.
If $G$ is a graph, then $N_{G}(v)=\{u\in V(G):uv\in E(G)\}$ is the neighborhood of the vertex $v$ in $G$.
The symbols $\omega(G)$ and $\alpha(G)$ denote, respectively, the clique number and the independence number of $G$.
A graph $Sp\neq K_{0}$ is said to be split if there exists a clique $K$ in $Sp$ and an independent set $S$ in $Sp$ such that
$V(Sp)=K\dot{\cup}S.$
In such a case, we say that the pair $(K,S)$ is a s-partition for $Sp$. Directly from the definition, it follows that the property of being split is hereditary. We call clique vertices the elements of $K$, and independent vertices the elements of $S$. Two s-partitions $(K,S)$ and $(K^{\prime},S^{\prime})$ for $Sp$ are equal if $K=K^{\prime}$ and $S=S^{\prime}$.
@@
If $Sp$ is a split graph, the notation $Sp(K,S)$ means that we are referring to $Sp$ together with a s-partition $(K,S)$ for it. Unless otherwise specified, we implicitly assume that both $K$ and $S$ are nonempty, and the adjacency matrix of $Sp$ has the form
@@
after applying the ordering $(K,S)$ to $V(Sp)$. Therefore, the blocks JIJ-I and 0 of $Sp$ are square, of size $|K|$ and $|S|$ respectively, while $R$, the bipartite adjacency matrix associated to $Sp$, is of size $|K|\times|S|$.
@@ 3 Null space of split graphs
In the analysis of split graphs, the structure of the adjacency matrix often leads us to consider the column space of the submatrix $R$ and its relationship with the all-ones vector $\operatorname{1\!\!1}$. The next lemma establishes a general matrix-theoretic criterion that is pivotal for such analysis. It characterizes the membership of a vector in the image of a matrix via the containment of nullspaces, a formulation that proves especially useful when applied to rank-one matrices like $J=\operatorname{1\!\!1}\operatorname{1\!\!1}^{t}$.
@@
The nullspace of a graph $G$, denoted by $\operatorname{\mathcal{N}}(G)$, is the nullspace of its adjacency matrix. If $G$ is a graph, then the support of $G$, denoted by $\operatorname{Supp}(G)$, is the set of all vertices $v$ of $G$ such that $x_{v}\neq 0$ for some vector $x\in\operatorname{\mathcal{N}}(G)$, see [jaume2018null]. Consequently, $V(G)-\operatorname{Supp}(G)$ is the set of all vertices $v$ of $G$ such that $x_{v}=0$ for all $x\in\operatorname{\mathcal{N}}(G)$. We say that $G$ has an independent support if $\operatorname{Supp}(G)$ is an independent set of $G$.
@@
Let
$A(Sp)=\begin{pmatrix}J-I&R\\ R^{t}&0\end{pmatrix}$
be the adjacency matrix of a split graph $Sp$ with bipartition $(K,S)$. Then, we have the following:
@@
    From the hypothesis and part (1), we have $Rx_{S}=(I-J)x_{K}=x_{K}-Jx_{K}$. Since $\operatorname{1\!\!1}^{t}x_{K}=\operatorname{1\!\!1}^{t}x_{K}=0$, we get $Jx_{K}=0$, so $Rx_{S}=x_{K}$. Therefore, $x_{K}\in\operatorname{\mathcal{R}}(R)$. Also from part (1), $R^{t}x_{K}=0$, so $x_{K}\in\operatorname{\mathcal{N}}(R^{t})$. For any real matrix $R$, we have $\operatorname{\mathcal{R}}(R)\perp\operatorname{\mathcal{N}}(R^{t})$, hence $x_{K}=0$.
@@
    By Lemma 3.1 with $A=R$ and $B=J=\operatorname{1\!\!1}\operatorname{1\!\!1}^{t}$, we have $\operatorname{\mathcal{N}}(R^{t})\subset\operatorname{\mathcal{N}}(J)$. Now, for any $(x_{K},x_{S})\in\operatorname{\mathcal{N}}(Sp)$, by (1) we have $R^{t}x_{K}=0$, so $x_{K}\in\operatorname{\mathcal{N}}(R^{t})\subset\operatorname{\mathcal{N}}(J)$, meaning $\operatorname{1\!\!1}^{t}x_{K}=0$. Then by part (4), $x_{K}=0$. Therefore, $\operatorname{Supp}(Sp)\subset S$.∎
@@
It is important to note that the converse of item (5) of Theorem 3.2 is not true. In other words, there exist split graphs with $\operatorname{Supp}(S)\subset S$ and $\operatorname{1\!\!1}\notin\operatorname{\mathcal{R}}(R)$. To see that, consider the split graph in Figure 2.
@@
Figure 2: Split graph $Sp(\{1,2,3,4\},\{5,6,7,8,9\})$ with $\operatorname{Supp}(Sp)\subset S$, but $\operatorname{1\!\!1}\notin\operatorname{\mathcal{R}}(R)$.
Direct computations shows that $\operatorname{Supp}(S)=\{5,\ldots,9\}=S$. However, the equation $Rx=\operatorname{1\!\!1}$ has no solution because
@@
Let $Sp(K,S)$ be a split graph with bipartite adjacency matrix $R$. Suppose there exists a vertex $v\in K$ such that its neighborhood in $S$ can be partitioned as
@@
for some subset $W\subset K$. Then for any $y\in\operatorname{\mathcal{R}}(R)$, we have
@@
Since $y\in\operatorname{\mathcal{R}}(R)$, there exists $x\in\mathbb{R}^{|S|}$ such that $Rx=y$. Recall that for any vertex $w\in K$, the entry $y_{w}=(Rx)_{w}=R_{w*}x$ equals the sum of $x_{i}$ over all neighbors $i\in N_{Sp}(w)\cap S$.
@@
$y_{v}\!=\!R_{v*}x\!=\!\sum_{i\in N_{Sp}(v)\cap S}x_{i}\!=\!\sum_{u\in W}\Big(\sum_{i\in N_{Sp}(u)\cap S}x_{i}\Big)\!=\!\sum_{u\in W}R_{u*}x\!=\!\sum_{u\in W}y_{u}.\qed$
The former proposition shows that the image of $R$ must satisfy certain linear constraints determined by the combinatorial structure of the split graph. In particular, if a vertex’s neighborhood is the disjoint union of other vertices’ neighborhoods, then their corresponding coordinates in any vector $y\in\operatorname{\mathcal{R}}(R)$ are linearly dependent.
Theorem 3.4.
@@
be the adjacency matrix of a split graph $Sp$ with bipartition $(K,S)$ such that $|K|\geq 2$. The clique-kernel of $Sp$ is defined as
$\operatorname{Cker}{\left(Sp\right)}:=\operatorname{\mathcal{N}}(R^{t})\cap\operatorname{\mathcal{R}}((I-J)^{-1}R)\subset\mathbb{R}^{|K|}.$
@@
The clique-kernel $\operatorname{Cker}{\left(Sp\right)}$ captures precisely the part of the nullspace that involves the clique vertices. When this subspace is trivial, all support is confined to the independent set $S$.
The relationship between the sizes of the clique and independent set in a split graph has implications for its singularity. The following result establishes simple combinatorial criteria that guarantee when a split graph must be singular based solely on the cardinalities of its bipartition. In particular, we show that an imbalance in favor of the independent set always forces singularity, while in the balanced case, the singularity of the entire graph reduces to the singularity of the bipartite adjacency matrix. This provides an bridge between the graph’s global spectral properties and the local structure encoded in the matrix $R$.
@@
    ($\Rightarrow$) If $Sp$ is singular, then $\operatorname{nl}(Sp)>0$. By Theorem 3.2(3), we have $\operatorname{nl}(Sp)\geq\operatorname{nl}(R)$. If $\operatorname{nl}(R)=0$, then any $(x_{K},x_{S})\in\operatorname{\mathcal{N}}(Sp)$ must satisfy $x_{K}=0$ (otherwise $x_{K}$ would be a nonzero vector in $\operatorname{\mathcal{N}}(R^{t})$, but $\operatorname{\mathcal{N}}(R^{t})=0$ since $\operatorname{nl}(R)=0$ and $R$ is a square matrix). Then from $(I-J)x_{K}=Rx_{S}$ we get $Rx_{S}=0$, so $x_{S}\in\operatorname{\mathcal{N}}(R)=0$. Thus $(x_{K},x_{S})=0$, contradicting $\operatorname{nl}(Sp)>0$. Therefore $\operatorname{nl}(R)>0$, so $R$ is singular.
@@
    The condition $\deg_{Sp}(x)=\deg_{Sp}(y)$ for all $x,y\in K$ implies that all vertices in $K$ have the same number of neighbors in $S$, since within $K$ they are all mutually adjacent (each has degree |K|1|K|-1 within $K$). Thus, there exists $d\geq 0$ such that $|N_{Sp}(x)\cap S|=d$ for all $x\in K$.

### v2

@@ Abstract
We study the nullspace of the adjacency matrix of split graphs, whose vertex set can be partitioned into a clique and an independent set. We introduce the clique-kernel, a subspace that decides whether clique vertices lie in the support of a kernel eigenvector, and we prove that its dimension is at most one. This yields the the formula
$\operatorname{nl}(Sp)=\operatorname{nl}(R)+\dim(\operatorname{Cker}{\left(Sp\right)}),$
which fully describes the nullity of a split graph in terms of the biadjacency submatrix $R$. We also analyze unbalanced split graphs through the concept of swing vertices and characterize the structure of their kernel supports. Furthermore, we study the behavior of the nullspace under Tyshkevich composition and derive a closed formula for the determinant. These results provide a unified algebraic framework for understanding when a split graph is singular and how its combinatorial structure determines its nullspace.
@@ 1 Introduction
The study of spectral properties of graphs, particularly the nullspace of the adjacency matrix, has emerged as a fundamental area in algebraic graph theory with significant implications for both theoretical mathematics and applied sciences. A graph $G$ is said to be singular if its adjacency matrix $A(G)$ is singular, that is, if the dimension of the nullspace of $A(G)$ is at least one. The nullity $\operatorname{nl}(G)$ of a graph $G$ is the multiplicity of zero as an eigenvalue of $A(G)$ or, equivalently, the dimension of the nullspace of $A(G)$. Understanding the structure of singular graphs and characterizing their nullspace has been a subject of intensive investigation since 1957 when von Collatz and Sinogowitz posed the problem, see [20].
@@
In her seminal work [16] Sciriha introduced the notion of core vertices (those vertices whose entries in a kernel eigenvector are nonzero) and laid the foundations for the study of singular graphs. This approach led to a systematic characterization of singular graphs in terms of minimal configurations and core structures. Subsequently, numerous researchers have contributed to this framework, including studies on nut graphs (core graphs of nullity one) [15], maximal core sizes [17], and the relationship between core and Fiedler vertices [1]. Jaume and Molina, see [11], presented a decomposition of trees in terms of the supported vertices, which turn out to be equivalent to the classical Edmond-Gallai decomposition, see [12].
@@
Graphs whose vertex set can be partitioned into a clique and an independent set were first singled out by Földes and Hammer in 1977, (see [4] and [5]), under the name split graphs. The same authors had earlier studied them implicitly as the graphs that do not contain an induced copy of $2K_{2}$, $C_{4}$ or $C_{5}$ ([7]). The term “split” was coined to stress that the vertex set is split into two sets whose induced subgraphs are as dissimilar as possible. Subsequent monographs of Golumbic [8] and Mahadev–Peled [13] embedded the class into the wider hierarchy of perfect graphs and made the notion standard in graph theory. Many NP-hard problems that are intractable for general graphs admit efficient algorithms when restricted to split graphs [9]. Moreover, split graphs can be recognized in linear time, and classical problems such as finding maximum cliques and maximum independent sets can be solved efficiently on this class [6].
@@
The spectral properties of split graphs have received considerable attention. Threshold graphs, which form a proper subclass of split graphs characterized by the absence of induced $P_{4}$, $C_{4}$, and $2K_{2}$ [13], have been extensively studied from a spectral perspective. Sciriha [18] investigated the spectrum of threshold graphs and established that singular threshold graphs possess kernel eigenvectors with at most two nonzero entries. Recent work by Panelo, Jaume, and Machado Toledo [14] on the core-nilpotent decomposition of threshold graphs further illuminates the algebraic structure of these important graph families.
@@
Despite the rich literature on both singular graphs and split graphs, a comprehensive study of the nullspace structure specifically for split graphs has been lacking. The block structure of the adjacency matrix of a split graph $Sp$,
$A(Sp)=\begin{pmatrix}J-I&R\\ R^{t}&0\end{pmatrix}$,
@@
where $J$ is the all-ones matrix, $I$ is the identity matrix, and $R$ encodes adjacencies between the clique $K$ and independent set $S$, naturally suggests a systematic approach to understanding the nullspace of $A(Sp)$. This structure allows us to decompose the nullspace problem into more manageable components and to establish clear relationships between the rank properties of $R$ and the nullity of $Sp$.
@@
    A detailed analysis of unbalanced split graphs (Section 4), i.e., those whose s-partition is not unique. We prove that the nullspace structure depends critically on whether the set of swing vertices (those which can belong to either part of an s-partition) forms a clique or an independent set.
@@
    For split graphs with supported clique vertices (Section 5), we prove that the nullity equals $\operatorname{nl}(R)+1$ when the support intersects the clique non-trivially, and we provide explicit formulas relating kernel eigenvectors to the adjugate matrix when $\operatorname{nl}(Sp)=1$ (Section 6).
@@
    Applications to the Tyshkevich composition (Section 7), demonstrating how our nullspace characterizations extend to composite structures.
@@
    An explicit determinant formula for split graphs (Section 8) expressed in terms of the matrix $R$ using the Schur complement and the matrix determinant lemma, which facilitates computational verification of singularity.
Our results extend and complement existing work on singular graphs by exploiting the special structure of split graphs. While Sciriha’s general theory [16] applies to arbitrary graphs, the block decomposition of split graph adjacency matrices enables us to obtain more precise characterizations. Furthermore, our treatment of unbalanced split graphs and swing vertices reveals phenomena that have not been systematically explored in the literature on singular graphs.
The remainder of this paper is organized as follows. Section 2 introduces notation and provides background on split graphs. Section 3 establishes general facts about the kernel of split graphs, including fundamental relationships between the nullspaces of $R$, $R^{t}$ and $A(Sp)$. Section 4 analyzes the support of unbalanced split graphs and its relation with swing vertices. Section 5 focuses on split graphs whose support intersects the clique, and Section 6 specializes in those with nullity 1. Section 7 is about the nullspace of a Tyshkevich composition. Section 8 derives an explicit determinant formula.
Throughout this work, we demonstrate that the nullspace of split graphs exhibits rich structural properties that can be understood through the interplay between the clique-independent set s-partition and the rank properties of the biadjacency submatrix $R$. Our results provide both theoretical insights into the spectral theory of structured graph classes and computational tools for determining singularity and nullity of split graphs.
@@ 2 Notation
If $A$ is a matrix, then $A_{ij}$ denotes the $(i,j)$-entry of $A$ (i.e., the entry of $A$ in the intersection of row $i$ and column $j$), $A_{i*}$ denotes the row $i$ of $A$, and $A_{*j}$ denotes the column $j$ of $A$. If $A$ is square, we denote the adjugate of $A$ by $\operatorname{adj}(A)$. The symbol $e_{i}$ refers to the $i$-th canonical vector, i.e., the vector with a “1” in the $i$-th position and zeros elsewhere. By $I$ we denote the identity matrix. Sometimes, we write a singleton set $\{x\}$ simply as “xx”.
We denote by $A(G)$ the adjacency matrix of $G$. The range of $A$, i.e., the vector space spanned by the columns of $A$, is denoted by $\operatorname{\mathcal{R}}(A)$. Its dimension, the rank of $A$, is denoted by $\operatorname{rk}(A)$. If $G$ is a graph, then $\operatorname{\mathcal{R}}(G)$ and $\operatorname{rk}(G)$ denote, respectively, the range and the rank of its adjacency matrix. The nullspace of $A$ is denoted by $\operatorname{\mathcal{N}}(A)$. Its dimension, the nullity of $A$, is denoted by $\operatorname{nl}(A)$. If $G$ is a graph, then $\operatorname{\mathcal{N}}(G)$ and $\operatorname{nl}(G)$ denote, respectively, the nullspace and the nullity of its adjacency matrix.
Let $U$ be a set of vectors of some vector space $V$; by $\langle U\rangle$ we denote the subspace spanned by all the linear combinations of vectors in $U$. We just write $\langle u\rangle$ instead of $\langle\{u\}\rangle$. $J$ is the “all-ones” matrix. Since $\operatorname{rk}(J)=1$, $\operatorname{nl}(J)$ is the number of columns of $J$ minus 1. $\operatorname{1\!\!1}=(1,\ldots,1)^{t}$ is the “all-ones” vector (i.e., the matrix $J$ with just one column), where “t” means transpose. If $A$ is a nonzero matrix with non-negative entries, note that $\operatorname{1\!\!1}\notin\operatorname{\mathcal{N}}(A)$. Clearly, $\operatorname{\mathcal{R}}(J)=\langle\operatorname{1\!\!1}\rangle$. If $x=(x_{1},\ldots,x_{n})^{t}\in\mathbb{R}^{n}$, we have $\operatorname{1\!\!1}^{t}x=\sum_{i=1}^{n}x_{i}$. Clearly, $\operatorname{1\!\!1}^{t}x=0$ if and only if $x\in\operatorname{\mathcal{N}}(J)$.
If $G$ is a graph, then $N_{G}(v)=\{u\in V(G):uv\in E(G)\}$ is the neighborhood of the vertex $v$ in $G$ (we can write $N(v)$ if $G$ is clear from the context). The symbols $\omega(G)$ and $\alpha(G)$ denote, respectively, the clique number and the independence number of $G$.
A graph $G\neq(\varnothing,\varnothing)$ is said to be split if there exists a clique $K$ in $G$ and an independent set $S$ in $G$ such that $V(G)=K\dot{\cup}S$. In such a case, we use the notation $Sp$ instead of $G$ and we say that the pair $(K,S)$ is an s-partition for $Sp$. Directly from the definition, it follows that every induced subgraph of a split graph is also split. We call clique vertices the elements of $K$, and independent vertices the elements of $S$. Two s-partitions $(K,S)$ and $(K^{\prime},S^{\prime})$ for $Sp$ are equal if $K=K^{\prime}$ and $S=S^{\prime}$.
@@
If $Sp$ is a split graph, the notation $Sp(K,S)$ means that we are referring to $Sp$ together with an s-partition $(K,S)$ for it. Unless otherwise specified, we implicitly assume that both $K$ and $S$ are nonempty, and the adjacency matrix of $Sp$ has the form
@@
after applying the ordering $(K,S)$ to $V(Sp)$. Therefore, the blocks JIJ-I and 0 of $Sp$ are square, of order $|K|$ and $|S|$ respectively, while $R$, the biadjacency (i.e., bipartite-adjacency) matrix associated to $Sp$, is of size $|K|\times|S|$.
@@ 3 Nullspace of split graphs
When studying split graphs one is naturally led to investigate the column space of the biadjacency matrix $R$ and its interplay with the all-ones vector $\operatorname{1\!\!1}$. The next lemma establishes a general matrix-theoretic criterion that is pivotal for such analysis. It characterizes the membership of a vector in the image of a matrix via the containment of nullspaces, a formulation that proves especially useful when applied to rank-one matrices like $J=\operatorname{1\!\!1}\operatorname{1\!\!1}^{t}$.
@@
Let $A$ be any matrix with $n$ columns. The support of $A$, denoted by $\operatorname{Supp}(A)$, is the set of all $v\in[n]=\{1,\ldots,n\}$ such that $x_{v}\neq 0$ for some vector $x\in\operatorname{\mathcal{N}}(A)$. If $G$ is a graph, then the support of $G$, denoted by $\operatorname{Supp}(G)$, is the support of its adjacency matrix, i.e., set of all vertices $v$ of $G$ such that $x_{v}\neq 0$ for some vector $x\in\operatorname{\mathcal{N}}(G)$ (see [11]). Consequently, $V(G)\setminus\operatorname{Supp}(G)$ is the set of all vertices $v$ of $G$ such that $x_{v}=0$ for all $x\in\operatorname{\mathcal{N}}(G)$.
@@
If $Sp(K,S)$ is a split graph, then we have the following.
@@
    From the hypothesis and part (1), we have $Rx_{S}=(I-J)x_{K}=x_{K}-Jx_{K}$. Since $\operatorname{1\!\!1}^{t}x_{K}=0$, we get $Jx_{K}=0$, so $Rx_{S}=x_{K}$. Therefore, $x_{K}\in\operatorname{\mathcal{R}}(R)$. Also from part (1), $R^{t}x_{K}=0$, so $x_{K}\in\operatorname{\mathcal{N}}(R^{t})$. Hence, $x_{K}=0$, because $\operatorname{\mathcal{R}}(R)\cap\operatorname{\mathcal{N}}(R^{t})=\{0\}$ for any real matrix $R$.
@@
    By Lemma 3.1, we have $\operatorname{\mathcal{N}}(R^{t})\subset\operatorname{\mathcal{N}}(J)$. Now, for any $(x_{K},x_{S})\in\operatorname{\mathcal{N}}(Sp)$, by (1) we have $R^{t}x_{K}=0$, so $x_{K}\in\operatorname{\mathcal{N}}(R^{t})\subset\operatorname{\mathcal{N}}(J)$, meaning $\operatorname{1\!\!1}^{t}x_{K}=0$. Then by part (4), $x_{K}=0$. Therefore, $\operatorname{Supp}(Sp)\subset S$.∎
@@
It is important to note that the converse of item (5) of Theorem 3.2 is not true. In other words, there exist split graphs with $\operatorname{Supp}(S)\subset S$ and $\operatorname{1\!\!1}\notin\operatorname{\mathcal{R}}(R)$. To see that, consider the split graph $Sp(K,S)$ in Figure 2.
@@
Figure 2: Split graph $Sp(K,S)$ with $\operatorname{Supp}(Sp)\subset S$, but $\operatorname{1\!\!1}\notin\operatorname{\mathcal{R}}(R)$.
Direct computations shows that $\operatorname{Supp}(Sp)=\{5,\ldots,9\}=S$. However, the equation $Rx=\operatorname{1\!\!1}$ has no solution because
@@
Let $Sp(K,S)$ be a split graph with a vertex $v\in K$ such that
@@
for some subset $W\subset K$. Then, for any $y\in\operatorname{\mathcal{R}}(R)$, we have
@@
Since $y\in\operatorname{\mathcal{R}}(R)$, there exists $x\in\mathbb{R}^{|S|}$ such that $Rx=y$. Recall that for any vertex $w\in K$, the entry $y_{w}=(Rx)_{w}=R_{w*}x$ equals the sum of $x_{i}$ over all neighbors $i\in N(w)\cap S$.
@@
$y_{v}\!=\!R_{v*}x\!=\!\sum_{i\in N(v)\cap S}x_{i}\!=\!\sum_{u\in W}\Big(\sum_{i\in N(u)\cap S}x_{i}\Big)\!=\!\sum_{u\in W}R_{u*}x\!=\!\sum_{u\in W}y_{u}.\qed$
The former proposition shows that the image of $R$ must satisfy certain linear constraints determined by the combinatorial structure of the split graph. In particular, if a vertex’s neighborhood is the disjoint union of other neighborhoods, then their corresponding coordinates in any vector $y\in\operatorname{\mathcal{R}}(R)$ are linearly dependent.
@@
be the adjacency matrix of a split graph $Sp(K,S)$ such that $|K|\geq 2$. The clique-kernel of $Sp$ is defined as
$\operatorname{Cker}{\left(Sp\right)}=\operatorname{\mathcal{N}}(R^{t})\cap\operatorname{\mathcal{R}}((I-J)^{-1}R)\subset\mathbb{R}^{|K|}.$
Theorem 3.4.
Let $Sp(K,S)$ be a split graph such that $|K|\geq 2$.
@@
The clique-kernel of $Sp$ captures precisely the part of the nullspace that involves the clique vertices. When this subspace is trivial, all the supported vertices of $Sp$ are independent.
The relationship between the sizes of the clique and independent set in a split graph has implications for its singularity. The following result establishes simple combinatorial criteria that guarantee when a split graph must be singular based solely on the cardinalities of its s-partition. In particular, we show that an imbalance in favor of the independent set always forces singularity, while in the balanced case, the singularity of the entire graph reduces to the singularity of the biadjacency matrix. This provides a bridge between the graph’s global spectral properties and the local structure encoded in the matrix $R$.
@@
    ($\Rightarrow$) If $Sp$ is singular, then $\operatorname{nl}(Sp)>0$. By Theorem 3.2(3), we have $\operatorname{nl}(Sp)\geq\operatorname{nl}(R)$. If $\operatorname{nl}(R)=0$, then any $(x_{K},x_{S})\in\operatorname{\mathcal{N}}(Sp)$ must satisfy $x_{K}=0$ (otherwise $x_{K}$ would be a nonzero vector in $\operatorname{\mathcal{N}}(R^{t})$, but $\operatorname{\mathcal{N}}(R^{t})=\{0\}$ since $\operatorname{nl}(R)=0$ and $R$ is a square matrix). Then from $(I-J)x_{K}=Rx_{S}$ we get $Rx_{S}=0$, so $x_{S}\in\operatorname{\mathcal{N}}(R)=\{0\}$. Thus $(x_{K},x_{S})=0$, contradicting $\operatorname{nl}(Sp)>0$. Therefore $\operatorname{nl}(R)>0$, so $R$ is singular.
@@
    The condition $\deg(x)=\deg(y)$ for all $x,y\in K$ implies that all vertices in $K$ have the same number of neighbors in $S$, since within $K$ they are all mutually adjacent (each has degree |K|1|K|-1 within $K$). Thus, there exists $d\geq 0$ such that $|N(x)\cap S|=d$ for all $x\in K$.